OpenStudy (astrophysics):

Is this sequence convergent or divergent? I don't really see either nor monotone.

OpenStudy (astrophysics):

\[a_n = \cos(n \pi/2)\] math questions account xD @ganeshie8

OpenStudy (astrophysics):

It's bounded between -1 and 1, so it's not monotone...

OpenStudy (astrophysics):

But is it convergent then, because of the bounds, mhm.

ganeshie8 (ganeshie8):

since the limit is DNE the sequence is divergent

OpenStudy (astrophysics):

OH

ganeshie8 (ganeshie8):



OpenStudy (astrophysics):

Yeah i just drew out the graph as well lol, to make sure what the heck is going on

ganeshie8 (ganeshie8):

it flip flops between 1 and -1 i guess

OpenStudy (astrophysics):

I remember reading another one similar to this I think it was \[a_n =(-1)^n\] so this one would also flip flop right? So therefore it's divergent as it doesn't exist either.

OpenStudy (astrophysics):

\[\left\{ -1,1,-1,1,-1... \right\}\]